OpenStudy (anonymous):

Write each expression as a single radical and then simplify your answer. a) √4√2x b) √3√10a^7b

OpenStudy (anonymous):

I'm just trying to clarify, is it \[\sqrt{4}\sqrt{2x}\] and \[\sqrt{3}\sqrt{10a ^{7b}}\] ?

OpenStudy (anonymous):

It is square root ^4√2x and same with the second one as well. If you could please help me understand this @jazzc